OpenStudy (anonymous):

Jenny plans to have two children but doesn't know if they are boy-boy, girl-girl, or boy-girl. What is the probability that she will have two boys?

OpenStudy (anonymous):

0.5 0.25 0.75 1.00

OpenStudy (anonymous):

How are you supposed to find that out?

OpenStudy (anonymous):

that's the question

OpenStudy (anonymous):

1/2 * 1/2 = 1/4

OpenStudy (anonymous):

Took him that long? :/

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@mikeperkins1997 Do you know the answer now? I know it because of him.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

1/4 = 0.25

OpenStudy (anonymous):

Exactly.